# Bounds and detection of MeV-scale dark matter annihilation to neutrinos

### v1

@@ Abstract
Current and most upcoming neutrino detectors can only reach a dark matter annihilation cross section to neutrinos larger than the standard freeze-out value, but they open intriguing detection avenues for non-standard dark matter paradigms. An important corollary of such non-standard scenarios is relic dark matter annihilation after neutrino decoupling, an effect that does not modify the neutrino temperature and was neglected in constraining the standard MeV-scale dark matter masses. By combining the contributions from entropy injection during neutrino decoupling and nonthermal neutrino energy release after decoupling, we derive significant constraints on the annihilation cross section to neutrinos, which in some mass regimes turn out to be stronger than the current bounds. Furthermore, we find that the lower bounds on dark matter masses become inconclusive under the recent data releases from the DESI, SPT-3G, and ACT collaborations. These bounds determine the extent to which upcoming neutrino detectors will probe dark matter annihilation into neutrinos.
@@ I Introduction
MeV-scale DM annihilating into neutrinos is a promising target for several low-energy neutrino experiments, including Borexino Borexino:2010zht ; Borexino:2019wln , KamLAND KamLAND:2011bnd ; KamLAND:2021gvi , Super-Kamiokande (SK) Super-Kamiokande:2005wtt ; Super-Kamiokande:2008ecj ; Super-Kamiokande:2010tar ; Super-Kamiokande:2011lwo ; Olivares-DelCampo:2017feq , Hyper-Kamiokande (HK) Olivares-DelCampo:2018pdl ; Bell:2020rkw , and JUNO JUNO:2015zny ; Akita:2022lit ; JUNO:2023vyz . These detectors can search for neutrino fluxes originating from galactic/extragalactic sources, such as DM annihilation and diffuse supernova neutrino background. However, the current, and even the proposed, sensitivities of these experiments can hardly reach the annihilation cross section required to produce the correct DM relic abundance, $\langle\sigma v\rangle_{\rm th}\simeq 3\times 10^{-26}\text{cm}^{3}/\text{s}$  Arguelles:2019ouk ; Dutta:2022wdi , as predicted by the standard thermal DM freeze-out paradigm Kolb:1990vq , and hence cannot decisively test the standard scenario. Nevertheless, this limitation can be turned into an advantage, as these experiments provide invaluable avenues for probing more general DM scenarios, where the annihilation cross section into neutrinos is larger while still consistent with the observed relic density. In fact, neutrino detection may become the only viable way to uncover the origin and particle nature of DM when it interacts with SM particles solely via the neutrino portal Batell:2017cmf ; Ballett:2019pyw ; Blennow:2019fhy ; Li:2022bpp ; Xu:2023xva .
@@
Depending on the datasets used, the resulting DM mass bounds can significantly impact the extent to which upcoming neutrino detectors, such as JUNO with 20 years of data, can probe MeV-scale DM annihilation into neutrinos. We illustrate this by highlighting the interplay between the detection sensitivity to $\langle\sigma v\rangle$ and modifications to $N_{\rm eff}$ in the range of 0.20.20.20.20.40.40.40.4, which is favored by several studies aiming to alleviate the Hubble tension DiValentino:2021izs ; Schoneberg:2021qvd ; Kamionkowski:2022pkx ; DiValentino:2025sru . We also find that the constraints arising from relic DM annihilation are comparable to current bounds obtained from observations of astrophysical neutrino fluxes. For DM masses approaching the lower limits, these cosmological constraints even surpass those from the current detectors.
@@ II Thermal MeV DM with relic annihilation
Constraints from neutrino detectors shown in shaded regions of Fig. 1 include the Borexino collaboration Borexino:2010zht ; Borexino:2019wln where upper limits of neutrino flux was used in Ref. Arguelles:2019ouk , the updated measurements from KamLAND where the $J$-factor of the angular-averaged intensity over the whole Milky Way is taken to be 1.3 KamLAND:2021gvi (KamLAND 2022), data of three different SK phases Super-Kamiokande:2005wtt ; Super-Kamiokande:2008ecj ; Super-Kamiokande:2010tar ; Super-Kamiokande:2011lwo used in Ref. Olivares-DelCampo:2017feq (SK 2018), and data from SK phase IV Super-Kamiokande:2013ufi performed in Ref. Arguelles:2019ouk (SK 2019). The forecast sensitivity from JUNO 2016 was performed in Ref. Arguelles:2019ouk based on the background estimates for diffuse supernova background searches JUNO:2015zny , while that from JUNO 20yrs was performed in Ref. Akita:2022lit with 20 years of data-taking under the standard Navarro-Frenk-White DM profile.
We see from Fig. 1 that the varying lower mass bounds significantly affects the extent to which the upcoming neutrino detection can probe MeV-scale thermal DM with an annihilation cross section larger than $\langle\sigma v\rangle_{\rm th}$. As seen in Fig. 1, with the severe bounds from SK Olivares-DelCampo:2017feq ; Arguelles:2019ouk , there is still large parameter space that will be targeted by JUNO for $m_{\rm DM}\lesssim 15$ MeV. Note that there is also large parameter space that can be probed with the forecast HK sensitivity Horiuchi:2008jz ; Bell:2020rkw for $10^{-25}\text{cm}^{3}/\text{s}\lesssim\langle\sigma v\rangle<10^{-24}~{}\text{cm}^{3}/\text{s}$, but it works for $m_{\rm DM}>30$ MeV and hence is not shown here.
@@ IV Conclusion
The latest measurements of $N_{\rm eff}$ after Planck 2018 have a significant impact on thermal MeV-scale DM that has a large annihilation cross section into neutrinos. They also affect the feasibility of probing such DM particles by observing extra neutrino flux in neutrino detectors. Applying the updated data releases from DESI, SPT-3G and ACT, we have found that the lower mass bound for MeV-scale DM becomes inconclusive, where significant parameter space can still be targeted by upcoming neutrino experiments. In doing this, we have included the relic DM annihilation effect on late-time contributions to $N_{\rm eff}$. This treatment strengthens the upper bounds of the annihilation cross section to neutrinos, which are comparable with or stronger than the current bounds from neutrino detectors. We have also illustrated the complementary probes of real scalar DM, by showing that an excess of $N_{\rm eff}$ at $0.2-0.4$ remarkably resides in the region that can be probed by upcoming JUNO experiments.

### v2

@@ Abstract
Current and most upcoming neutrino detectors can only reach a dark matter annihilation cross section to neutrinos larger than the standard freeze-out value, but they open intriguing detection avenues for non-standard dark matter paradigms. An important corollary of these non-standard scenarios is relic dark matter annihilation after neutrino decoupling, which was previously overlooked in constraining MeV-scale dark matter. However, by combining the contributions from entropy injection during neutrino decoupling and from nonthermal neutrino energy release after decoupling, we derive significant constraints on the annihilation cross section to neutrinos, which in some mass regimes become stronger than the current bounds. Furthermore, we find that the lower bounds on dark matter masses become inconclusive under the recent data releases from the DESI, SPT-3G, and ACT collaborations. These bounds determine the extent to which upcoming neutrino detectors will probe dark matter annihilation into neutrinos.
@@ I Introduction
MeV-scale DM annihilating into neutrinos is a promising target for several low-energy neutrino experiments, including Borexino Borexino:2010zht ; Borexino:2019wln , KamLAND KamLAND:2011bnd ; KamLAND:2021gvi , Super-Kamiokande (SK) Super-Kamiokande:2005wtt ; Super-Kamiokande:2008ecj ; Super-Kamiokande:2010tar ; Super-Kamiokande:2011lwo ; Olivares-DelCampo:2017feq , Hyper-Kamiokande (HK) Olivares-DelCampo:2018pdl ; Bell:2020rkw , and JUNO JUNO:2015zny ; Akita:2022lit ; JUNO:2023vyz . These detectors can search for neutrino fluxes originating from galactic/extragalactic sources, such as DM annihilation and diffuse supernova neutrino background. However, the current, and even the proposed, sensitivities of these experiments can hardly reach the annihilation cross section required to produce the correct DM relic abundance, $\langle\sigma v\rangle_{\rm th}\simeq 3\times 10^{-26}\text{cm}^{3}/\text{s}$  Steigman:2012nb ; Arguelles:2019ouk ; Dutta:2022wdi ; Chu:2023jyb , as predicted by the standard thermal DM freeze-out paradigm Kolb:1990vq , and hence cannot decisively test the standard scenario. Nevertheless, this limitation can be turned into an advantage, as these experiments provide invaluable avenues for probing more general DM scenarios, where the annihilation cross section into neutrinos is larger while still consistent with the observed relic density. In fact, neutrino detection may become the only viable way to uncover the origin and particle nature of DM when it interacts with SM particles solely via the neutrino portal Batell:2017cmf ; Ballett:2019pyw ; Blennow:2019fhy ; Li:2022bpp ; Xu:2023xva .
@@
Depending on the datasets used, the resulting DM mass bounds can significantly impact the extent to which upcoming neutrino detectors, such as JUNO with 20 years of data, can probe MeV-scale DM annihilation into neutrinos. We illustrate this by highlighting the interplay between the detection sensitivity to $\langle\sigma v\rangle$ and modifications to $N_{\rm eff}$ in the range of 0.20.20.40.4, which is favored by several studies aiming to alleviate the Hubble tension DiValentino:2021izs ; Schoneberg:2021qvd ; Kamionkowski:2022pkx ; DiValentino:2025sru . We also find that the constraints arising from relic DM annihilation are comparable to current bounds obtained from observations of astrophysical neutrino fluxes. For DM masses approaching the lower limits, these cosmological constraints even surpass those from the current detectors.
@@ II Thermal MeV DM with relic annihilation
Constraints from neutrino detectors shown in shaded regions of Fig. 1 include the Borexino collaboration Borexino:2010zht ; Borexino:2019wln where upper limits of neutrino flux was used in Ref. Arguelles:2019ouk , the updated measurements from KamLAND where the $J$-factor of the angular-averaged intensity over the whole Milky Way is taken to be 1.3 KamLAND:2021gvi (KamLAND 2022), data of three different SK phases Super-Kamiokande:2005wtt ; Super-Kamiokande:2008ecj ; Super-Kamiokande:2010tar ; Super-Kamiokande:2011lwo used in Ref. Olivares-DelCampo:2017feq (SK 2018), and data from SK phase IV Super-Kamiokande:2013ufi performed in Ref. Arguelles:2019ouk (SK 2019). The forecast sensitivity from JUNO 2016 was performed in Ref. Arguelles:2019ouk based on the background estimates for diffuse supernova background searches JUNO:2015zny , while that from JUNO 20yrs was performed in Ref. Akita:2022lit with 20 years of data-taking under the standard Navarro-Frenk-White DM profile. The standard thermal WIMP case shown in the black dotted line corresponds to $\langle\sigma v\rangle_{\rm th}\simeq 3\times 10^{-26}\text{cm}^{3}/\text{s}$. See Refs. Steigman:2012nb ; Chu:2023jyb for more precise determination of $\langle\sigma v\rangle_{\rm th}$.
We see from Fig. 1 that the varying lower mass bounds significantly affects the extent to which the upcoming neutrino detection can probe MeV-scale thermal DM with an annihilation cross section larger than $\langle\sigma v\rangle_{\rm th}$. As seen in Fig. 1, with the severe bounds from SK Olivares-DelCampo:2017feq ; Arguelles:2019ouk , there is still large parameter space that will be targeted by JUNO for $m_{\rm DM}\lesssim 15$ MeV.
While not explicitly shown in Fig. 1, it is worth mentioning that there is also large parameter space that can be probed with the forecast HK sensitivity Horiuchi:2008jz ; Hyper-Kamiokande:2018ofw ; Bell:2020rkw for $10^{-25}\text{cm}^{3}/\text{s}\lesssim\langle\sigma v\rangle<10^{-24}~\text{cm}^{3}/\text{s}$ and $m_{\rm DM}>10$ MeV. DM with mass above 10 MeV up to 1 GeV may also be covered by future tonne-scale DM direct detection experiments such as DARWIN and ARGO McKeen:2018pbb , though they are under planning phases. With the increased sensitivity from DUNE DUNE:2015lol ; Capozzi:2018dat , a DM mass above 100 MeV and the annihilation cross section down to the level of $10^{-24}~\text{cm}^{3}/\text{s}$ can be reached Arguelles:2019ouk .
@@ IV Conclusion
The latest measurements of $N_{\rm eff}$ after Planck 2018 have a significant impact on thermal MeV-scale DM that has a large annihilation cross section into neutrinos. They also affect the feasibility of probing such DM particles by observing extra neutrino flux in neutrino detectors. Applying the updated data releases from DESI, SPT-3G and ACT, we have found that the lower mass bound for MeV-scale DM becomes inconclusive, where significant parameter space can still be targeted by upcoming neutrino experiments. In doing this, we have included the relic DM annihilation effect on late-time contributions to $N_{\rm eff}$. This treatment strengthens the upper bounds of the annihilation cross section to neutrinos, which are comparable with or stronger than the current bounds from neutrino detectors. We have also illustrated the complementary probes of real scalar DM, by showing that an excess of $N_{\rm eff}$ at 0.20.40.2-0.4 remarkably resides in the region that can be probed by upcoming JUNO experiments.